# SP-VIO: Robust and Efficient Filter-Based Visual Inertial Odometry with State Transformation Model and Pose-Only Visual Description

### v1

@@ IV Measurement Model Based on PO Theory
IV-A PO Description of Multiple View Geometry
@@
$\boldsymbol{p}_{{C}_{i}}=\frac{1}{{}^{{C}_{i}}{Z}_{f}}{{}^{{C}_{i}}{p}_{f}}=\frac{1}{{Z}_{f}^{{C}_{i}}}{{}_{w}^{{C}_{i}}\boldsymbol{R}}(^{G}\boldsymbol{p}_{f}-^{G}\boldsymbol{t}_{{C}_{i}})$ (25)
@@
IV-B Measurement Model
In MSCKF, the reprojection error is related to the camera poses and 3D features, but the system state does not contain 3D features. Therefore, it is necessary to use the left nullspace to restore the standard EKF format before filtering updates.
@@ IV-B Measurement Model
V DST-RTS Correction Strategy
@@
Due to the sensitivity of cameras, visual tracking failure is easy to occur in challenging environments, resulting in discontinuous visual observation and affecting the localization accuracy of VINS.
@@
In this section, we first analyze the coping strategy and output trajectory of the mainstream VIO algorithm under visual deprived conditions, then propose a correction strategy based on RTS backtracking smoothing [30], and analyze the advantages of this method over existing works.
@@ V-A Mainstream VIO Algorithms
During visual deprivation, MSCKF [10] performs inertial navigation solution and updates the error covariance matrix. As the linearization error of inertial navigation accumulates, the error covariance increases gradually. When the observation is recovered, MSCKF will correct the estimated state to the direction of visual measurement, resulting in a jump in the estimated result, as shown in Fig. 4. If the visual measurement is more accurate, the jump direction will reduce the current error to some extent.
@@
From Fig. 5, it can be found that the cumulative error during visual deprivation is included in the subsequent navigation results, which is related to the nonlinear optimization function constructed by VINS-Mono [3]:
$\mathop{min}\limits_{\mathcal{X}}=\left\{||{\boldsymbol{r}_{p}}-{\boldsymbol{H}_{p}}{\mathcal{X}}||^{2}+\sum_{k\in\mathcal{B}}||{\boldsymbol{r}_{k}}||^{2}+\sum_{j\in\mathcal{C}}||{\boldsymbol{r}_{j}}||^{2}\right\}$ (35)
where ${\boldsymbol{r}_{p}}-{\boldsymbol{H}_{p}}{\mathcal{X}}$ is the residual of prior information after marginalization, $\mathcal{B}$ and $\mathcal{C}$ are the IMU measurements set and visual measurements set in the sliding window respectively, ${\boldsymbol{r}_{k}}$ and ${\boldsymbol{r}_{j}}$ are the residual terms of IMU measurements and visual measurements respectively.
As shown in (35), Since the construction of the nonlinear optimization function is mainly based on the relative relationship between IMU measurements and visual measurements, the estimated results can only correct the local relative error, and cannot affect the cumulative error during the visual deprivation. Also for this reason, VINS-Mono does not show the estimated jump in Fig. 4.
@@ V-B DST-RTS Backtracking Smoothing
Combining the DST-EKF proposed in Section III, this subsection provides a correction strategy basd on RTS smoother, which can be used without loop closure. Specific steps are as follows:
@@
The above steps are based on DST-EKF, where (1), (2) remarked as ’DST-EKF-forward’, and (3) remarked as ’DST-EKF-RTS’. The timing chart of the proposed method is shown in Fig. 7:
@@
Considering that the global frame $\{G\}$ before and after the visual interruption may be inconsistent, we take the velocity of body frame ${{}^{b}{\boldsymbol{v}}}$ as the measurement, and construct the measurement model as follows:
@@
$\delta\boldsymbol{x}_{v}$ $=\begin{bmatrix}\delta\boldsymbol{x}_{b}^{T}&\delta\boldsymbol{s}^{T}\end{bmatrix}^{T}$ (37)
where $\delta\boldsymbol{x}_{b}$ is consistent with (5), $\delta\boldsymbol{s}$ is the scale deviation of VIO after visual recovery , $\boldsymbol{n}_{v}$ is the measurement noise, and $\boldsymbol{Z}_{v}$ can be defined as:
@@
$={{}^{b}{\boldsymbol{\widetilde{v}}_{IMU}}}-{{}^{b}{\boldsymbol{v}}_{VIO}}-{{}^{b}{\boldsymbol{\widetilde{v}}}_{VIO}}{\delta\boldsymbol{s}}$ (38)
@@
where ${{}^{b}{\boldsymbol{\widetilde{v}}_{IMU}}}$ is the velocity estimated by the INS solution, and ${{}^{b}{\boldsymbol{\widetilde{v}}}_{VIO}}$ is the velocity provided by the VIO system.
As shown in (V-B), the observation jacobian matrix $\boldsymbol{H}_{v}$ can be represented as:
@@
In summary, compared with the mainstream algorithms mentioned in Section V-A, our proposed correction strategy can effectively correct the accumulated errors during visual deprivation and does not depend on the loop closure.
@@ VI Experimental Evaluation
This section provides a comprehensive performance evaluation of our proposed SP-VIO and compares it with the current SOTA VIO algorithms, such as OpenVINS [6] and VINS-Mono [3]. The benchmark datasets include both open-source and personal datasets, and all contain time-synchronized IMU data, images, and ground truth, where a detailed overview is provided in Table I. To quantitatively evaluate the localization accuracy of the relevant algorithms, we calculate absolute translation error (ATE) [71] across the full experimental trajectory by the EVO tools [72]. All experiments are conducted on a standard laptop (Intel Core i7-10875H CPU @ 2.80GHZ). Since the RTS relocalization algorithm is mainly used for discontinuous observation conditions, we dedicate a subsection to evaluating it.
@@
VI-A Ablation experiments on SP-VIO
To verify the validity of the new system model and observational model, as well as the performance of complete SP-VIO, we adopt MSCKF as the baseline system to construct three algorithms: SP-VIO(Without ST), SP-VIO(Without PO), and SP-VIO, respectively, and carried out ablation experiments on the Kitti Odometry dataset [33].
This dataset includes different acquisition scenarios, such as urban, rural, and highway, with complex motion conditions and long distances, which is conducive to comparing the performance of VIO algorithms. In order to ensure the fairness of this experiment, we initialize the four algorithms by assigning values, and provide the same preprocessing features to avoid the influence of other variables on the experimental results. Additionally, the imu data of ’00’ was lost in some periods, and ’03’ has been removed from the official website, so we chose another 9 sequences for the evaluation.
TABLE II: RMSE in Kitti Odometry Dataset in Meters
Sequence Distance MSCKF SP-VIO(Without ST) SP-VIO(Without PO) SP-VIO
01 $2457.57$ $27.54$ $20.06/{27.16\%}^{1}$ $20.48/25.64\%$ $\boldsymbol{19.27}/\boldsymbol{30.02\%}$
02 $2685.71$ $14.35$ $12.79/10.87\%$ $13.01/9.34\%$ $\boldsymbol{12.01}/\boldsymbol{16.30\%}$
[…]
05 $2208.69$ $16.23$ $16.19/0.25\%$ $15.95/1.72\%$ $\boldsymbol{15.57}/\boldsymbol{4.07\%}$
06 $1219.59$ $5.94$ $2.78/53.2\%$ $5.72/3.7\%$ $\boldsymbol{2.7}/\boldsymbol{54.54\%}$
07 $692.83$ $3.56$ $3.02/15.17\%$ $2.82/20.79\%$ $\boldsymbol{2.8}/\boldsymbol{21.35\%}$
08 $3226.71$ $19.34$ $9.17/52.59\%$ $13.97/27.77\%$ $\boldsymbol{6.77}/\boldsymbol{64.99\%}$
09 $1710.34$ $13.08$ $7.55/42.28\%$ $9.38/28.29\%$ $\boldsymbol{6.67}/\boldsymbol{49.01\%}$
10 $921.22$ $8.10$ $6.32/21.98\%$ $6.16/23.95\%$ $\boldsymbol{6.05}/\boldsymbol{25.31\%}$
Average $1724$ $12.09$ $8.68/25.84\%$ $9.77/18.31\%$ $\boldsymbol{8.01}/\boldsymbol{33.75\%}$
  • 1
    The percentage of localization error decrease in this algorithm is compared to MSCKF.

### v2

@@ III-B Measurement Model Based on PO Visual description
III-B1 PO Description of Multiple View Geometry
@@
$\boldsymbol{p}_{{C}_{i}}=\frac{1}{{}^{{C}_{i}}{Z}_{f}}{{}^{{C}_{i}}{p}_{f}}=\frac{1}{{Z}_{f}^{{C}_{i}}}{{}_{G}^{{C}_{i}}\boldsymbol{R}}(^{G}\boldsymbol{p}_{f}-^{G}\boldsymbol{t}_{{C}_{i}})$ (18)
@@
III-B2 PO-Based State Update
In MSCKF-based VIO, the reprojection error is related to the camera poses and 3D features, but the system state does not contain 3D features. Therefore, it is necessary to use the left nullspace to restore the standard EKF format before filtering updates.
@@ III-B2 PO-Based State Update
IV Observability Analysis
Observability is an important metric for consistency analysis of filter-based VIO systems, and the observability matrix for the linearized system can be construct following [20, 19]:
$\boldsymbol{O}=\begin{bmatrix}\boldsymbol{O}_{0}\\ \boldsymbol{O}_{1}\\ \vdots\\ \boldsymbol{O}_{k}\end{bmatrix}=\begin{bmatrix}\boldsymbol{H}_{0}\\ \boldsymbol{H}_{1}{\boldsymbol{\Phi}_{1,0}}\\ \vdots\\ \boldsymbol{H}_{k}{\boldsymbol{\Phi}_{k,0}}\end{bmatrix}$ (28)
In this section, we start with the unobservable subspace of filter-based VIO systems, first perform observability analysis of several traditional multi-view geometry based VIO systems, and then discuss the effect of the introduction of pose-only visual description on the system’s observability and whether there is consistency improvement.
IV-A Multi-View geometry Based VIO Systems
Due to the measurement model format of MSCKF-based VIO[7, 10, 9, 11, 6], observability analysis is difficult to perform directly. It is proved in [8] that the observability analysis of EKF-SLAM and MSCKF is equivalent. Therefore, the unobservable subspace of several algorithms are calculated on the basis of the linearized model of EKF-SLAM in this subsection.
Assuming that a single visual feature is observed at time $k$, the state vector is defined as:
${\boldsymbol{x}}=\begin{bmatrix}{}_{{b}_{k}}^{G}\boldsymbol{q}^{T}&{}^{G}\boldsymbol{v}_{{b}_{k}}^{T}&{}^{G}\boldsymbol{p}_{{b}_{k}}^{T}&\boldsymbol{b}_{{g}_{k}}^{T}&\boldsymbol{b}_{{a}_{k}}^{T}&{}^{G}\boldsymbol{p}_{f}^{T}\end{bmatrix}^{T}$ (30)
From [20], the unobservable subspace of Std-EKF based VIO systems [7, 6] can be expressed as:
${\boldsymbol{N}_{k}^{Std}}=\begin{bmatrix}{\boldsymbol{0}_{3\times 3}}&{\boldsymbol{g}^{G}}\\ {\boldsymbol{0}_{3\times 3}}&-[{{}^{G}{\boldsymbol{v}}_{b_{0}}}\times]{\boldsymbol{g}^{G}}\\ {\boldsymbol{I}_{3\times 3}}&-[{{}^{G}{\boldsymbol{p}}_{b_{0}}}\times]{\boldsymbol{g}^{G}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 3}}\\ {\boldsymbol{I}_{3\times 3}}&[{{}^{G}{\boldsymbol{p}}_{f}}\times]{\boldsymbol{g}^{G}}\end{bmatrix}$ (31)
From [20], the unobservable subspace in (31) is susceptible to system linearization change, resulting in spurious information inflation to the unobservable direction causing inconsistency problems.
To solve the inconsistent problem, [61] proposed the right invariant (RL)-VINS, which adopts right invariant errors for VINS states except for IMU biases. The unobservable subspace of RL-VINS can be computed as[69]:
${\boldsymbol{N}_{k}^{RI}}=\begin{bmatrix}{\boldsymbol{0}_{3\times 3}}&{\boldsymbol{g}^{G}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{I}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 3}}\\ {\boldsymbol{I}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\end{bmatrix}$ (32)
As shown in (32), the null space of RI-VINS is unrelated to the state vector and can directly satisfy the unobservable properties of VINS without Jacobian modifications. However, the disadvantage of RI-VINS is that the visual features participate in state covariance propagation in Lie-Group representation. Assuming that the state vector contains $m$ features, the computation cost of covariance propagation for Std-ekf based VIO systems is only around $\textbf{O}(m)$ but can rise to $\textbf{O}(m^{2})$ for RI-VINS[69], which significantly increases the computational cost.
Different from RI-VINS, DST-EKF based VIO only carries out state transformation for velocity and position errors without modifying the visual features state representation so that the unobservable subspace can be expressed as:
${\boldsymbol{N}_{k}^{DST}}=\begin{bmatrix}{\boldsymbol{0}_{3\times 3}}&{\boldsymbol{g}^{G}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{I}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 3}}\\ {\boldsymbol{I}_{3\times 3}}&-[{{}^{G}{\boldsymbol{p}}_{f}}\times]{\boldsymbol{g}^{G}}\end{bmatrix}$ (33)
From (31) to (33), DST-EKF based VIO has better consistency than Std-ekf based VIO, and does not suffer from the computational costs associated with visual feature amplification compared to RI-VINS. However, the feature estimation will affect its unobservable subspace,, and the system will suffer from the inconsistency caused by the feature linearization error when the feature estimation is not accurate.
In addition, similar work has been done with Dri-VINS[69], which reduces the computational cost by decoupling features from Lie group representations based on RI-VINS. Dri-VINS and DST-EKF based VIO have similar observability but are also plagued by feature linearization errors. Therefore, FEJ method [17] is introduced in [69] to maintain the system’s consistency.
Since multi-view geometry based VIO is associated with 3D features, when 3D features are not included in the state vectors, the linearization error introduced by feature triangulation will destroy the system’s consistency, but adding 3D features to the state covariance propagation will lead to a surge in computing costs and limit the application of the system under different conditions.
IV-B Pose-Only Visual description Based VIO Systems
Unlike multi-view geometry based VIO, PO based VIO is decoupled from 3D features, so 3D features are not added to the state vector in observability analysis. Assuming that a single visual feature is observed in the images at three moments, and the images at time $\alpha$ and time $\beta$ are taken as the left/right-base views, the state vectors of PO-EKF and SP-VIO at time $k$ can be defined as:
$\boldsymbol{x}$ $=\begin{bmatrix}\boldsymbol{x}_{b}^{T}&\boldsymbol{x}_{c}^{T}\end{bmatrix}^{T}$ (34)
$\boldsymbol{x}_{b}$ $=\begin{bmatrix}{}_{{b}_{k}}^{G}\boldsymbol{q}^{T}&{}^{G}\boldsymbol{v}_{{b}_{k}}^{T}&{}^{G}\boldsymbol{p}_{{b}_{k}}^{T}&\boldsymbol{b}_{{g}_{k}}^{T}&\boldsymbol{b}_{{a}_{k}}^{T}\end{bmatrix}^{T}$
$\boldsymbol{x}_{c}$ $=\begin{bmatrix}{}_{{b}_{\alpha}}^{G}\boldsymbol{q}^{T}&{}^{G}\boldsymbol{p}_{{b}_{\alpha}}^{T}&{}_{{b}_{\beta}}^{G}\boldsymbol{q}^{T}&{}^{G}\boldsymbol{p}_{{b}_{\beta}}^{T}\end{bmatrix}^{T}$
From (25) and (34), the Jacobian matrix of PO-EKF and SP-VIO at time $k$ is defined as:
[…]
Combine [20] and (35), the unobservable subspace of PO-EKF can be computed as:
${\boldsymbol{N}_{k}^{PO}}=\begin{bmatrix}{\boldsymbol{0}_{3\times 3}}&\left({{}^{G}\boldsymbol{p}_{b_{k}}}-{{}^{G}\boldsymbol{p}_{b_{\alpha}}}-{{}^{G}\boldsymbol{v}_{b_{\alpha}}}+{\boldsymbol{g}^{G}}\right)\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{I}_{3\times 3}}&-{[\left({{}^{G}\boldsymbol{p}_{b_{k}}}-{{}^{G}\boldsymbol{p}_{b_{\alpha}}}\right)\times]}\\ &\left({{}^{G}\boldsymbol{p}_{b_{k}}}-{{}^{G}\boldsymbol{p}_{b_{\alpha}}}-{{}^{G}\boldsymbol{v}_{b_{\alpha}}}+{\boldsymbol{g}^{G}}\right)\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&\left({{}^{G}\boldsymbol{p}_{b_{k}}}-{{}^{G}\boldsymbol{p}_{b_{\alpha}}}-{{}^{G}\boldsymbol{v}_{b_{\alpha}}}+{\boldsymbol{g}^{G}}\right)\\ {\boldsymbol{I}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{I}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\end{bmatrix}$ (36)
Similar to (31), although the unobservable subspace of PO-EKF is decouples from the visual features, it is also susceptible to linearization errors, resulting in inconsistency problems.
After replacing Std-EKF with DST-EKF, the unobservable subspace of SP-VIO can be expressed as:
${\boldsymbol{N}_{k}^{SP}}=\begin{bmatrix}{\boldsymbol{0}_{3\times 3}}&{\boldsymbol{g}^{G}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{I}_{3\times 3}}&-{[\left({{}^{G}\boldsymbol{p}_{b_{k}}}-{{}^{G}\boldsymbol{p}_{b_{\alpha}}}\right)\times]}{\boldsymbol{g}^{G}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{g}^{G}}\\ {\boldsymbol{I}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{0}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\\ {\boldsymbol{I}_{3\times 3}}&{\boldsymbol{0}_{3\times 1}}\end{bmatrix}$ (37)
As shown in (37), the unobservable subspace of SP-VIO is less affected by system linearization changes than PO-EKF. Meanwhile, compared with the multi-view geometry based VIO systems mentioned in Section IV-A, SP-VIO can achieve better consistency while maintaining higher computational efficiency after decoupling from 3D visual features.
V Enhanced SP-VIO Combined with DST-RTS Correction Strategy
@@
Due to camera’s sensitivity, visual tracking failure is easy to occur in challenging environments, which results in discontinuous visual measurements and affects VINS’s localization accuracy.
@@
In this section, we first analyze the coping strategy and output trajectory of the mainstream VIO algorithm under visual deprived conditions, and then propose an enhanced RTS correction strategy [30] combined with the DST-EKF proposed in Section V to improve the robustness of SP-VIO under visual deprivation conditions.
@@ V-A Mainstream VIO Algorithms
During visual deprivation, MSCKF [6] performs an inertial navigation solution and updates the error covariance matrix. As the linearization error of inertial navigation accumulates, the error covariance increases gradually. When the observation is recovered, MSCKF will correct the estimated state to the direction of visual measurement, resulting in a jump in the estimated result, as shown in Fig. 4. If the visual measurement is more accurate, the jump direction will somewhat reduce the current error.
@@
Since the construction of the nonlinear optimization function in VINS-Mono [3] is mainly based on the relative relationship between IMU and visual measurements, the estimated results can only correct the local relative error and cannot affect the cumulative error during the visual deprivation, as shown in Fig. 5. Also for this reason, VINS-Mono does not show the estimated jump in Fig. 4.
@@ V-B DST-RTS Backtracking Smoothing
As shown in [70], the propagation process of linearized systems with matrix Lie group element will be independent of the state estimate, so the RTS smoothing method with Lie group representation can better deal with the motion trajectory on poor initial state estimation. From Section IV, DST-EKF has similar propagation characteristics with the filter algorithms represented by matrix Lie group. Therefore, DST-EKF is adopted in this section to reconstruct RTS smoothing to deal with the motion trajectory under visual deprivation condition.
The specific steps to enhanced RTS backtracking smoothing are as follows:
@@
Notably, (1), (2) remarked as ’DST-EKF-forward’, and (3) remarked as ’DST-EKF-RTS’. The timing chart of the proposed method is shown in Fig. 7:
@@
Considering that there is a scale deviation in the system’s estimation after visual recovery, we add the scale error $\delta\boldsymbol{s}$ to the error state vector of (5):
$\delta\boldsymbol{x}_{b}$ $=\begin{bmatrix}{}_{{b}_{k}}^{G}\boldsymbol{\phi}^{T}&\delta^{G}\boldsymbol{v}_{{b}_{k}}^{T}&\delta^{G}\boldsymbol{p}_{{b}_{k}}^{T}&\delta\boldsymbol{b}_{{g}_{k}}^{T}&\delta\boldsymbol{b}_{{a}_{k}}^{T}&{\delta\boldsymbol{s}}^{T}\end{bmatrix}^{T}$ (38)
Introducing the body frame’s velocity after visual recovery, the measurement model of DST-EKF-forward can be defined as:
@@
$={{}^{b}{\boldsymbol{\widetilde{v}}_{IMU}}}-{{}^{b}{\boldsymbol{v}}_{VIO}}-{{}^{b}{\boldsymbol{\widetilde{v}}}_{VIO}}{\delta\boldsymbol{s}}$
@@
where $\boldsymbol{n}_{v}$ is the measurement noise, ${{}^{b}{\boldsymbol{\widetilde{v}}_{IMU}}}$ is the velocity estimated by the INS solution, and ${{}^{b}{\boldsymbol{\widetilde{v}}}_{VIO}}$ is the velocity provided by the VIO system after visual recovery.
As shown in (39), the measurement jacobian matrix $\boldsymbol{H}_{v}$ can be represented as:
@@
In summary, combined with the enhanced DST-RTS smoothing strategy, our proposed SP-VIO can correct accumulated errors after visual recovery, and does not depend on the loop closure.
@@ VI Experimental Evaluation
This section provides a comprehensive performance evaluation of our proposed SP-VIO, including the ablation experiments and comparisons with the current SOTA VIO algorithms, such as OpenVINS [7] and VINS-Mono [3]. Real-world experiments include both open-source and personal datasets, and all contain time-synchronized IMU data, images, and ground truth, where a detailed overview is provided in Table I. To quantitatively evaluate the localization accuracy of the relevant algorithms, we calculate absolute translation error (ATE) [71] across the entire experimental trajectory using the EVO tools [72]. All experiments are conducted on a standard laptop (Intel Core i7-10875H CPU @ 2.80GHZ). Since the RTS relocalization algorithm is mainly used for discontinuous observation conditions, we dedicate a subsection to evaluating it.
@@
VI-A Ablation experiments
To verify our proposed algorithm, ablation experiments were first performed on the simulation trajectory and KITTI Odometry dataset [33], and the comparison algorithms included: Std-EKF(baseline system)[7], DST-EKF(modified by Section III-A), and PO-EKF(modified by Section III-B).
VI-A1 Monte-Carlo Simulations
The simulation trajectory is shown in Fig. 8, corresponding to the three carriers of UAV, handheld device and commercial vehicle respectively, which can cover most motion situations in the real world environment. The inertial and camera measurements of the simulation are generated by [7], and the basic configuration is shown in Table II.
The overall results are shown in Table III and Fig. 9. From the results, SP-VIO showed the best localization performance on all simulation trajectory thanks to the reconstructed system model and measurement model, which allowed it to output more consistent navigation results under different motion states. In addition, DST-EKF and PO-EKF show suboptimal performance, but are not satisfactory on some sequences, which is consistent with the observability analysis results in Section IV.
(a) SIM_UAV
(b) SIM_Pedestrian
(c) SIM_CAR
Figure 8: Simulated trajectories. Left: SIM_UAV, middle: SIM_Pedestrian and right: SIM_CAR . Note that the red triangular denotes the start of the trajectory while the green square indicates the end.
TABLE II: Monte-Carlo Simulation Parameters
Parameter Value Parameter Value
IMU Freq. (hz) $400$ Cam Freq. (hz) $10$
Gyro White Noise $1.6968e^{-4}$ Gyro Random Walk $1.9393e^{-4}$
Accel. White Noise $2.0000e^{-3}$ Accel. White Noise $2.0000e^{-3}$
Max Cam Pts/Frame $100$ Max Cam Pts/Frame $40$
Max Clone Size $11$ Pt Feat. Rep. GLOBAL
Cam noise (pixel) $1$ Cam noise (pixel) false
TABLE III: RMSE for 50 Monte-Carlo Runs on 3 Simulated Trajectories in Meters
Sequence Distance Std-EKF DST-EKF PO-EKF SP-VIO
SIM_UAV $662$ $0.33$ $0.43$ $0.24$ $\boldsymbol{0.20}$
SIM_Pedestrian $293$ $0.48$ $0.14$ $0.41$ $\boldsymbol{0.10}$
SIM_CAR $3678$ $3.93$ $3.64$ $4.20$ $\boldsymbol{2.65}$
Average $1544$ $1.58$ $1.41$ $1.62$ $\boldsymbol{0.99}$
(a) SIM_UAV
(b) SIM_Pedestrian
(c) SIM_CAR
Figure 9: Average localization error for 50 Monte-Carlo runs of algorithms(Std-EKF, PO-EKF, DST-EKF, and SP-VIO) on different simulation trajectories, respectively.
VI-A2 Real-World Experiments
We further evaluate the SP-VIO with real-world KITTI benchmark, which includes different acquisition scenarios, such as urban, rural, and highway, with complex motion conditions and long distances. In order to ensure the fairness of this experiment, we initialize the four algorithms by assigning values, and provide the same preprocessing features to avoid the influence of other variables on the experimental results. Additionally, the imu data of ’00’ was lost in some periods, and ’03’ has been removed from the official website, so we chose another 9 sequences for the evaluation.
TABLE IV: RMSE in Kitti Odometry Dataset in Meters
Sequence Distance Std-EKF DST-EKF PO-EKF SP-VIO
01 $2457.57$ $27.54$ $20.06$ $20.48$ $\boldsymbol{19.27}$
02 $2685.71$ $14.35$ $12.79$ $13.01$ $\boldsymbol{12.01}$
04 $393.31$ $0.64$ $\boldsymbol{0.26}$ $0.5$ $\boldsymbol{0.26}$
05 $2208.69$ $16.23$ $16.19$ $15.95$ $\boldsymbol{15.57}$
06 $1219.59$ $5.94$ $2.78$ $5.72$ $\boldsymbol{2.7}$
07 $692.83$ $3.56$ $3.02$ $2.82$ $\boldsymbol{2.8}$
08 $3226.71$ $19.34$ $9.17$ $13.97$ $\boldsymbol{6.77}$
09 $1710.34$ $13.08$ $7.55$ $9.38$ $\boldsymbol{6.67}$
10 $921.22$ $8.10$ $6.32$ $6.16$ $\boldsymbol{6.05}$
Average $1724$ $12.09$ $8.68$ $9.77$ $\boldsymbol{8.01}$